# A Smooth Sea Never Made a Skilled SAILOR: Robust Imitation via Learning to Search

### v1

@@ 1 Introduction
We propose an algorithmic paradigm that allows us to acquire both of these capabilities without requiring any sources of human data beyond the standard imitation learning pipeline. In other words, a better recipe with the same ingredients. In particular, rather than merely learning a policy from expert demonstrations, we propose learning a local world model (WM) and a reward model (RM) from demonstration and base policy data (Ren et al., 2024b). By combining these components with a planning algorithm, we have the capability to, at test time, reason about how to recover from mistakes that the base policy makes by planning against our learned RM inside our learned WM. Thus, rather than mere imitation, we’re learning to search (L2S, Ratliff et al. (2009)) from expert demonstrations. Our key insight is that we can infer the latent search process required to recover from local mistakes from the same source of human data required for the standard behavioral cloning pipeline.
@@ 2 Related Work
After expending computation at test-time for a search procedure, it is often valuable to distill the search process back into the base policy, an approach known as expert iteration or dual policy iteration (Anthony et al., 2017; Sun et al., 2018). While such approaches have long demonstrated strong performance in continuous control (Levine and Koltun, 2013; Wang et al., 2025), they have attracted renewed interest in the context of LLM reasoning (Zelikman et al., 2022; Gandhi et al., 2024; Hosseini et al., 2024). SAILOR can be seen as a generalization of these ideas to continuous control problems with stochastic dynamics, visual observations, and unknown reward functions, which none of the above can handle directly. We ablate the value of ExIt-like updates and find they provide nontrivial improvement in policy performance. Recent work by Wu et al. (2025b) applies similar ideas on real-world visual manipulation problems but focuses on mode selection rather than more general behavior correction. SAILOR can be thought of fusing the paradigms of residual RL (Silver et al., 2018; Yuan et al., 2025; Ankile et al., 2024) and L2S by learning to search for residuals.
@@ 3.1 What is Learning to Search?
World Model. To avoid the complexity of modeling the dynamics of and planning directly in the space of high-dimensional observations, we adopt the Recurrent State-Space Model architecture (RSSM, Hafner et al. (2019, 2024)). This means our world model is composed of three key components, i.e. . The first, the encoder encodes the observation into latent space i.e. . The second, the latent dynamics model predicts the next latent after taking an action, i.e. . The third, the decoder tries to reconstruct the observation from the latent state, i.e. . In SAILOR, we train all three of these components on hybrid data, i.e. a mixture of and – see Appendix A for all loss functions. As argued by Ross and Bagnell (2012); Vemula et al. (2023); Ren et al. (2024b), while such a world model is only locally accurate on the expert’s and learner’s state distributions, a policy that looks good when evaluated in such a model is guaranteed to do well in the real world. 333While we do not investigate this in our experiments, the above theory still holds if the world model is trained on a wide data distribution that covers both the expert and learner’s visitation distribution. Thus, a particularly interesting future direction is to train a powerful foundation world model (Agarwal et al., 2025; Bruce et al., 2024; Parker-Holder et al., 2024) on internet-scale robotics datasets (Khazatsky et al., 2024; ONeill et al., 2024), before plugging it into the SAILOR framework to allow for wider-ranging recovery from mistakes.
@@ 4 Experiments
Benchmarks. We evaluate the efficacy of SAILOR on 12 challenging visual manipulation problems. This includes 3 tasks from Robomimic (Mandlekar et al., 2021b) {Lift, Can, Square}, 6 from RoboSuite (Zhu et al., 2020) {Door, Stack, Bread, Cereal, Round, BreadCanRandom}, and 3 from ManiSkill (Tao et al., 2025) {PullCube, PokeCube, LiftPeg}. These include diverse tasks: multi-step pick-and-place (BreadCanRandom), articulated object manipulation (Door), and tool use (Square, Round, PokeCube). For Robomimic tasks, we use the provided demonstrations, while we collect our own demonstrations using a SpaceMouse controller for the other suites. To solve each task, SAILOR is provided with a set of expert demonstrations , along with a budget specifying the maximum number of environment interactions it can perform. The observation space consists of RGB images from a wrist-mounted camera and a third-person camera mounted in front of the agent, and the proprioceptive states. More details are provided in App. C and D.
Algorithms. We compare three imitation learning methods: Diffusion Policies (DP, Chi et al. (2023)) trained via BC, a model-free inverse RL method (DPPO-IRL) that directly updates DP parameters via RL (DPPO, Ren et al. (2024a)) against a learned RM with a separate encoder, and SAILOR. For evaluation, we measure the Success Rate (SR) across rollouts, and report the mean and standard error obtained with 3 seeds. More details on the implementation of the methods are provided in App. A, B. All methods were trained on 1 NVIDIA 6000 Ada GPU with 48 GB of memory.
@@ 4.2 What Matters in Learning To Search?
Warm-Start. We found that allocating of the given interaction budget to a “warm-start phase” significantly boosts our final performance, as shown in the leftmost part of Fig. 9. We attribute this to the fact that having the WM, RM, and V accurate on ’s distribution reduces exploration in Phase II.
@@
Hybrid WM Updates. Another component that has a significant effect on model performance is using a mix of expert and learner data to update the WM. We observe in Fig. 9, center, that this “hybrid” fitting of the WM not only has the potential to improve initial model performance (as for PokeCube), but can also improve final performance (as for Cereal), reflecting the insights of Ren et al. (2024b).
@@
Expert Iteration. Finally, we find that the inclusion of the expert iteration subroutine boosted model performance, as shown in the rightmost part of Fig. 9. This can be attributed to the base policy knowing how to recover from mistakes its prior iterations would have made.
5 Conclusion
@@
In summary, we see that across a dozen visual manipulation tasks, SAILOR is able to train robust agents that recover from from the failures of the base policy via test-time planning and match the expert’s outcomes, all without any additional human data. Furthermore, scaling up the amount of data used to train a diffusion policy via behavioral cloning by 5-10 is unable to close this performance gap. We also see that our learned RM is able to detect nuanced failures, while the SAILOR agent is able to recover from them. As mentioned above, while for our experiments we use a base diffusion policy, MPPI planner, and Dreamer world model, the general SAILOR architecture is fully compatible with foundation behavior (e.g. VLAs) and world models trained on internet-scale data, which might unlock further levels of generalization and robustness – we leave this as a promising future direction. To close, we observe that in his oft-quoted essay The Bitter Lesson, Richard Sutton writes that:
@@ Acknowledgments
Appendix A Implementation Details
In this section we first describe the network architecture for each component of SAILOR in Sec. A.1 following details of training pipeline and hyperparameters in Sec. A.2
A.1 Network Architecture
@@ Appendix D Demonstrations
Appendix E Related Work (Extended)
Reward Models. Another approach to interactive imitation learning is inverse reinforcement learning (IRL, [Ng et al., 2000, Syed and Schapire, 2007, Ziebart et al., 2008, Ho and Ermon, 2016, Swamy et al., 2021]), which does not require human-in-the-loop queries. In IRL, one learns a classifier that maximally differentiates learner from expert behavior and uses it as a reward model (RM) for RL-based policy updates. Different forms of reward models include successor features [Jain et al., 2025], score matching [Wu et al., 2025a], and optimal transport metrics Haldar et al. [2023b, a]. Unfortunately, the RL step of IRL is often rather interaction-inefficient. Recent theoretical work has argued that rather than a global RL procedure, a local search procedure555By local search, we mean merely competing with the expert rather than the optimal policy for an adversarially chosen reward. For the former, there exist algorithms that avoid the worst-case, exponential-in-the-horizon exploration complexity of RL [Bagnell et al., 2003, Ross and Bagnell, 2014, Swamy et al., 2023]. is sufficient for performant imitation [Swamy et al., 2023, Espinosa-Dice et al., 2025]. SAILOR fits within the overarching algorithmic paradigm of local search IRL but learns to search rather than learning a policy directly.
World Models. World models (WMs, Ha and Schmidhuber [2018]) have been an integral component of impressive RL results in a variety of domains [Hafner et al., 2020, 2021, Jain et al., 2022, Hansen et al., 2022, Hafner et al., 2024, Hansen et al., 2024, Zhou et al., 2024, Bruce et al., 2024, Parker-Holder et al., 2024, Agarwal et al., 2025], including real-world robotics [Mendonca et al., 2021, Wu et al., 2023]. While our overall algorithmic framework is agnostic to the choice of WM architecture, we use Dreamer-style world models [Hafner et al., 2020] in our experiments due to their ubiquity. Popov et al. [2024] use rollouts in a learned world model to minimize a trajectory-level divergence between the expert and the learner at train time on autonomous driving problems. In contrast, we use the world model at test-time to enable recovery from just the mistakes the learner actually makes, which might be more computationally efficient than a global search procedure [Kearns et al., 2002].

### v2

@@ 1 Introduction
We propose an algorithmic paradigm that allows us to acquire both of these capabilities without requiring any sources of human data beyond the standard imitation learning pipeline. In other words, a better recipe with the same ingredients. In particular, rather than merely learning a policy from expert demonstrations, we propose learning a local world model (WM) and a reward model (RM) from demonstration and base policy data (Ren et al., 2024b). By combining these components with a planning algorithm, we have the capability to, at test time, reason about how to recover from mistakes that the base policy makes by planning against our learned RM inside our learned WM. Thus, rather than mere imitation, we’re learning to search (L2S, Ratliff et al. (2009)) from expert demonstrations. Our key insight is that we can infer the latent search process required to recover from local mistakes without any more human feedback (e.g., corrections) than a standard behavioral cloning pipeline.
@@ 2 Related Work
After expending computation at test-time for a search procedure, it is often valuable to distill the search process back into the base policy, an approach known as expert iteration or dual policy iteration (Anthony et al., 2017; Sun et al., 2018). While such approaches have long demonstrated strong performance in continuous control (Levine and Koltun, 2013; Wang et al., 2025), they have attracted renewed interest in the context of LLMs (Zelikman et al., 2022; Gandhi et al., 2024; Hosseini et al., 2024; Jain et al., 2025a). SAILOR can be seen as a generalization of these ideas to continuous control problems with stochastic dynamics, visual observations, and unknown reward functions, which none of the above can handle directly. We ablate the value of ExIt-like updates and find they provide nontrivial improvement in policy performance. Recent work by Wu et al. (2025b) applies similar ideas on real-world visual manipulation problems but focuses on mode selection rather than more general behavior correction. SAILOR can be thought of fusing the paradigms of residual RL (Silver et al., 2018; Yuan et al., 2025; Ankile et al., 2024) and L2S by learning to search for residuals.
@@ 3.1 What is Learning to Search?
World Model. To avoid the complexity of modeling the dynamics of and planning directly in the space of high-dimensional observations, we adopt the Recurrent State-Space Model architecture (RSSM, Hafner et al. (2019, 2024)). This means our world model is composed of three key components, i.e. . The first, the encoder encodes the observation into latent space i.e. . The second, the latent dynamics model predicts the next latent after taking an action, i.e. . The third, the decoder tries to reconstruct the observation from the latent state, i.e. . In SAILOR, we train all three of these components on hybrid data, i.e. a mixture of and – see Appendix B for all loss functions. As argued by Ross and Bagnell (2012); Vemula et al. (2023); Ren et al. (2024b), while such a world model is only locally accurate on the expert’s and learner’s state distributions, a policy that looks good when evaluated in such a model is guaranteed to do well in the real world. 333While we do not investigate this in our experiments, the above theory still holds if the world model is trained on a wide data distribution that covers both the expert and learner’s visitation distribution. Thus, a particularly interesting future direction is to train a powerful foundation world model (Agarwal et al., 2025; Bruce et al., 2024; Parker-Holder et al., 2024) on internet-scale robotics datasets (Khazatsky et al., 2024; ONeill et al., 2024), before plugging it into the SAILOR framework to allow for wider-ranging recovery from mistakes.
@@ 4 Experiments
Benchmarks. We evaluate the efficacy of SAILOR on 12 challenging visual manipulation problems. This includes 3 tasks from Robomimic (Mandlekar et al., 2021b) {Lift, Can, Square}, 6 from RoboSuite (Zhu et al., 2020) {Door, Stack, Bread, Cereal, Round, BreadCanRandom}, and 3 from ManiSkill (Tao et al., 2025) {PullCube, PokeCube, LiftPeg}. These include diverse tasks: multi-step pick-and-place (BreadCanRandom), articulated object manipulation (Door), and tool use (Square, Round, PokeCube). For Robomimic tasks, we use the provided demonstrations, while we collect our own demonstrations using a SpaceMouse controller for the other suites. To solve each task, SAILOR is provided with a set of expert demonstrations , along with a budget specifying the maximum number of environment interactions it can perform. The observation space consists of RGB images from a wrist-mounted camera and a third-person camera mounted in front of the agent, and the proprioceptive states. More details are provided in App. D and E.
Algorithms. We compare three imitation learning methods: Diffusion Policies (DP, Chi et al. (2023)) trained via BC, a model-free inverse RL method (DPPO-IRL) that directly updates DP parameters via RL (DPPO, Ren et al. (2024a)) against a learned RM with a separate encoder, and SAILOR. For evaluation, we measure the Success Rate (SR) across rollouts, and report the mean and standard error obtained with 3 seeds. More details on the implementation of the methods are provided in App. B, C. All methods were trained on 1 NVIDIA 6000 Ada GPU with 48 GB of memory.
@@ 4.2 What Matters in Learning To Search?
Warm-Start. We found that allocating of the interaction budget to a “warm-start phase” significantly boosts final performance, as shown in the leftmost part of Fig. 9. We attribute this to having the WM, RM, and V be accurate on ’s distribution, reducing exploration in Phase II. We find performance plateaus when using of the interaction budget for warm-starting (App. A.1).
@@
Hybrid WM Training. Another component that has a significant effect on model performance is using a mix of expert and learner data to update the WM. We observe in Fig. 9, center, that this “hybrid” fitting of the WM not only has the potential to improve initial model performance (as for PokeCube), but can also improve final performance (as for Cereal), reflecting the insights of Ren et al. (2024b).
@@
Expert Iteration. In Fig. 9, right, we see that the inclusion of the expert iteration subroutine boosts model performance. This can be attributed to the base policy knowing how to recover from mistakes its prior iterations would have made, effectively recycling test-time compute into a new base policy.
Residual Planning. We ablate the importance of using test-time planning to search for residuals online instead of a precomputed residual policy (RP). Similar to the residual planner, the RP takes as input both the latent state and base policy action. It is trained via standard actor-critic techniques as in Dreamer (Hafner et al., 2024). In Fig. 10, we see that across the board, explicit test-time planning out-performs a residual policy. As additional statistics, we also include numbers for the wall-clock time overhead of running MPPI, which we find to be somewhat limited. We also ablate the amount of MPPI iterations performed in the loop in App. A.2. There, we see fast convergence to an improved plan.
5 Discussion
@@
In summary, we see that across a dozen visual manipulation tasks, SAILOR is able to train robust agents that recover from from the failures of the base policy via test-time planning and match the expert’s outcomes, all without on-policy human feedback. Furthermore, scaling up the amount of data used to train a diffusion policy via behavioral cloning by 5-10 is unable to close this performance gap. We also see that our learned RM is able to detect nuanced failures, while the SAILOR agent is able to recover from them. As mentioned above, while for our experiments we use a base diffusion policy, MPPI planner, and Dreamer world model, the general SAILOR architecture is fully compatible with foundation behavior (e.g. VLAs) and world models trained on internet-scale data, which might unlock further levels of generalization and robustness – we leave this as a promising future direction.
Before we close, it is worth discussing two reasons why L2S is a particularly scalable algorithmic paradigm. The first is data-efficiency: as with any inverse RL approach that learns a verifier (i.e., reward model) from demonstrations rather than just a generator (i.e., policy), L2S requires less data than direct policy learning approaches on problems with a generation-verification gap (Swamy et al., 2025). Given we have far less data in robotics than in domains like language modeling, squeezing as much as we can out of every demonstration is of paramount importance. Furthermore, as we scale up task horizons, the size of the generation-verification gap grows. This means that L2S techniques could exhibit better “data scaling laws" with task complexity than behavioral cloning techniques.
The second reason is compute-efficiency: in L2S, we only need to plan at the states actually encountered at test-time, rather than all states we could potentially encounter as in a train-time procedure. This provides provable computational benefits (Kearns et al., 2002). Furthermore, expert iteration-style distillation lets us recycle this compute, so we never need to solve the same problem twice. Thus, L2S techniques could also exhibit better “compute scaling laws” than vanilla policy learning.
To close, we observe that in his oft-quoted essay The Bitter Lesson, Richard Sutton writes that:
@@ Acknowledgments
Appendix A Additional Results
We now discuss two additional ablations of the SAILOR stack.
A.1 Warm-Start Fraction
The warm-start fraction refers to the portion of total interaction budget we use to collect data with just the base policy (i.e., not performing any policy updates or using the planning stack). We then use this data to train the WM, RM and Critic. We performed an ablation of this fraction on the Square task with demonstrations.
In Fig. 11, we observe that using a warm-start fraction of 0.1 improves performance over not warm-starting. However, performance plateaus after using warm-start fraction of 0.2, which we used for all of our experiments.
A.2 Residual Search vs. Residual Policy
In Fig, 12, we compare the performance of SAILOR with different number of iterations of the MPPI planning process. We find that the performance converges after 2-3 iterations for most tasks and there is only a marginal improvement in later iterations, suggesting we could truncate SAILOR ’s search process early. Doing so adaptively on a state-wise basis is an interesting future research direction.
Additionally, we measured that SAILOR spends 0.014s time per iteration of MPPI, while it takes 0.066s for the base diffusion policy to generate actions, making test-time planning a somewhat limited overhead. This is likely because the MPPI planner operates on top of lower-dimensional latent states from the world model, rather than high-dimensional image observations like the base DP.
Lastly, we explored performing a final post-hoc distillation of the SAILOR agent into the base policy (Distilled DP) for fast inference at test-time. As seen in Fig. 12, the performance of Distilled DP is similar to the SAILOR agent on Lift and Can tasks. However, on the harder Square task, test-time search outperforms the Distilled DP. This is potentially due to the limited capacity of the diffusion policy for representing the full planning circuit for sufficiently complex tasks.
Appendix B Implementation Details
In this section we first describe the network architecture for each component of SAILOR in Sec. B.1 following details of training pipeline and hyperparameters in Sec. B.2
B.1 Network Architecture
@@ Appendix E Demonstrations
Appendix F Related Work (Extended)
Reward Models. Another approach to interactive imitation learning is inverse reinforcement learning (IRL, [Ng et al., 2000, Syed and Schapire, 2007, Ziebart et al., 2008, Ho and Ermon, 2016, Swamy et al., 2021]), which does not require human-in-the-loop queries. In IRL, one learns a classifier that maximally differentiates learner from expert behavior and uses it as a reward model (RM) for RL-based policy updates. Different forms of reward models include successor features [Jain et al., 2025b], score matching [Wu et al., 2025a], and optimal transport metrics Haldar et al. [2023b, a]. Unfortunately, the RL step of IRL is often rather interaction-inefficient. Recent theoretical work has argued that rather than a global RL procedure, a local search procedure555By local search, we mean merely competing with the expert rather than the optimal policy for an adversarially chosen reward. For the former, there exist algorithms that avoid the worst-case, exponential-in-the-horizon exploration complexity of RL [Bagnell et al., 2003, Ross and Bagnell, 2014, Swamy et al., 2023]. is sufficient for performant imitation [Swamy et al., 2023, Espinosa-Dice et al., 2025]. SAILOR fits within the overarching algorithmic paradigm of local search IRL but learns to search rather than learning a policy directly.
World Models. World models (WMs, Ha and Schmidhuber [2018]) have been an integral component of impressive RL results in a variety of domains [Hafner et al., 2020, 2021, Jain et al., 2022, Hansen et al., 2022, Hafner et al., 2024, Hansen et al., 2024, Zhou et al., 2025, Bruce et al., 2024, Parker-Holder et al., 2024, Agarwal et al., 2025], including real-world robotics [Mendonca et al., 2021, Wu et al., 2023]. While our overall algorithmic framework is agnostic to the choice of WM architecture, we use Dreamer-style world models [Hafner et al., 2020] in our experiments due to their ubiquity. Popov et al. [2024] use rollouts in a learned world model to minimize a trajectory-level divergence between the expert and the learner at train time on autonomous driving problems. In contrast, we use the world model at test-time to enable recovery from just the mistakes the learner actually makes, which might be more computationally efficient than a global search procedure [Kearns et al., 2002].